# Contrastive explainable clustering with differential privacy

### v1

@@ Appendix C Additional proofs for $k$-median algorithm
The solution of the $k$-median problem (with demands and a center fixed at a location $C$) can be formulated as finding the optimal solution of the following Integer program (IP):
@@
$y_{C}$ $=1\text{ for a fixed $C\in N$}$ (54)
$x_{CC}$ $=1\text{ for a fixed $C\in N$}.$ (55)
@@
Lemma 12.
(Lemma 4 of Charikar et al. [1999]) For any feasible integer solution $(x^{\prime},y^{\prime})$ for the input with modified demands, there is a feasible integer solution for the original input with cost at most $4\bar{C}_{LP}$ plus the cost of $(x^{\prime},y^{\prime})$ with demands $d^{\prime}$.
Lemma 13.
(Theorem 6 of Charikar et al. [1999]) There is a $1/2$-restricted solution $(x^{\prime},y^{\prime})$ of cost at most $2\bar{C}_{LP}$.
@@ Appendix D Additional proofs for $k$-means algorithm
(Full version of Lemma 5) Let $S$ be $1$-stable set and $O$ be the optimal set of $k$ centers, we have $\Delta(O)-3\Delta(S)+2R\geq 0$, where $R=\sum_{q\in P}\Delta(q,s_{o_{q}})$.
@@
(Full version of Lemma 6) With $R$ and $\alpha$ defined as above: $R\leq 2\Delta(O)+(1+2/\alpha)\Delta(S)$.
@@ D.1 Swap pairs mapping
Figure 2: A detailed visualization of our dataset (Charlottesville County, Virginia) and analysis includes (a) a Scatter plot of the full dataset with sampled points for contrastive analysis. (b) Comparison of $k$-median clustering with fixed and non-fixed centroids, both private and non-private. (c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-median with a fixed centroid. (d) Heatmap indicating $k$-median error costs based on fixed centroid choices.
@@ D.3.1 Datasets and Experimental Setup
Beyond the Albemarle County dataset discussed in the main body of the paper, we also conducted experiments on other datasets, including the Charlottesville city dataset and synthetic activity-based populations. Here, we provide a detailed analysis of these experiments.
Charlottesville City Dataset:
This dataset is part of the synthetic U.S. population, as described in  Chen et al. [2021], Barrett et al. [2009]. It comprises about 33K individuals and approximately 5600 activity locations visited by these individuals. The locations represent places where individuals perform various activities.
@@
The synthetic dataset was meticulously crafted to emulate the distribution and characteristics of real-world datasets, ensuring a balance between realism and controlled variability. Comprising 100 data points in a 2D space, this dataset was uniformly distributed, mirroring the range observed in the real datasets we analyzed.
The primary motivation behind this synthetic dataset is to provide a sandbox environment, free from the unpredictable noise and anomalies of real-world data. This controlled setting is pivotal in understanding the core effects of differential privacy mechanisms, isolating them from external confounding factors. The dataset serves as a foundational tool in our experiments, allowing us to draw comparisons and validate our methodologies before applying them to more complex, real-world scenarios.
@@ D.3.2 Experimental Results
For the Charlottesville city dataset, we observed similar trends to those in the Albemarle county dataset. As the $\epsilon$ value increased, prioritizing accuracy, there was a slight compromise in privacy. However, consistent with our hypothesis, the influence of the epsilon budget on the explainability of our outcomes was negligible.
Effect of Privacy on Explainability:
The contrastive explanations for the Charlottesville city dataset also showcased the resilience of our approach. Even with differential privacy perturbations, the quality of explanations remained consistent, emphasizing the robustness of our method.
Visualization:
Figure  2 - (a) showcases the Charlottesville city dataset. Blue dots represent the original dataset, while orange ’X’ markers indicate the sampled points for contrastives. The distribution of these points provides a visual representation of the data’s diversity and the regions we focused on for our contrastive analysis.
@@ D.3.3 Performance Evaluation:
For each $\epsilon$ value, we conducted 25 different runs on the Charlottesville city dataset. The average results were consistent with our findings from the Albemarle County dataset. It’s essential to note that these multiple invocations were solely for performance evaluation. In real-world applications, invoking private algorithms multiple times could degrade the privacy guarantee.
@@
Despite the distinct scales between the Charlottesville city and the Albemarle county datasets, we observed consistent patterns in the contrastive explanations. Specifically, as illustrated in Figure 2 - (b) and Figure  2 - (c), the contrastive explanations remained largely unaffected by variations in the $\epsilon$ budget. This consistency further reinforces our hypothesis that the epsilon budget has a negligible influence on the explainability of our outcomes, even when applied to datasets of different scales.
Impact of Fixed Centroids on Contrastive Error:
In Figure  2 - (d), we delve deeper into the effects of selecting different centroids as fixed centroids. It’s evident that the choice of a fixed centroid has a pronounced impact on the contrastive error. This underscores the importance of the centroid’s position concerning the data distribution. Some centroids, especially those closer to data-dense regions, tend to produce lower counterfactual errors, while others, particularly those near sparser regions or outliers, result in higher errors. This observation highlights the intricate relationship between centroid positioning and the resulting contrastive explanation’s accuracy.
@@ D.3.4 Conclusion:
The extended experiments on the Charlottesville city dataset further validate our approach’s efficacy. The balance between privacy and utility, the robustness of contrastive explanations, and the negligible impact of $\epsilon$ on explainability were consistent across datasets. These findings underscore the potential of our method for diverse real-world applications.

### v2

@@ Appendix C Additional proofs for $k$-median algorithm
The solution of the $k$-median problem (with demands and a center fixed at a location $z$ can be formulated as finding the optimal solution of the following Integer program (IP):
@@
$y_{z}$ $=1\text{ for a fixed }z\in N$ (44)
$x_{zz}$ $=1\text{ for a fixed }z\in N.$ (45)
@@ Appendix D Additional proofs for $k$-means algorithm
Let $S$ be $1$-stable set and $O$ be the optimal set of $k$ centers, we have $\Delta(O)-3\Delta(S)+2R\geq 0$, where $R=\sum_{q\in P}\Delta(q,s_{o_{q}})$.
@@
With $R$ and $\alpha$ defined as above: $R\leq 2\Delta(O)+(1+2/\alpha)\Delta(S)$.
@@ D.3.1 Datasets and Experimental Setup
Figure 3: A detailed visualization of our dataset (Charlottesville County, Virginia) and analysis includes (a) A scatter plot of the full dataset with 100 randomly selected points for contrastive analysis, chosen to provide a more comfortable and manageable subset for explanation purposes. (b) Comparison of $k$-median clustering with fixed and non-fixed centroids, both private and non-private. (c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-median with a fixed centroid.
Figure 4: The figure presents visualizations of the dataset (Heart Disease, UCI MLR) reduced to four dimensions (2, 4, 6, and 8), along with the corresponding analysis (a) A t-SNE scatter plot illustrating the high-dimensional data. (b) Comparison of $k$-means clustering with fixed and non-fixed centroids, both private and non-private. (c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-means with a fixed centroid.
Figure 5: The figure presents further visualizations of the dataset (Heart Disease, UCI MLR) introduced in the main paper, reduced to four dimensions (2, 4, 6, and 8), along with the corresponding analysis (a) A t-SNE scatter plot illustrating the high-dimensional data. (b) Comparison of $k$-median clustering with fixed and non-fixed centroids, both private and non-private. (c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-median with a fixed centroid.
Figure 6: The figure presents visualizations of the dataset (Breast Cancer - 30 features, UCI MLR) reduced to four dimensions (4, 6, 8, and 16), along with the corresponding analysis (a) A t-SNE scatter plot illustrating the high-dimensional data. (b) Comparison of $k$-means clustering with fixed and non-fixed centroids, both private and non-private. (c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-means with a fixed centroid.
In the main body of the paper, we focused on the Heart Disease dataset from the UCI Machine Learning Repository, which was reduced from 13 features to 8 features for visualization purposes. However, we have also conducted experiments on the same dataset with different dimensionality reductions to analyze the impact on our results. Furthermore, to demonstrate the robustness of our work, we have applied our analysis to additional datasets, including synthetic datasets and real-world datasets from Charlottesville, Albemarle and Breast Cancer Wisconsin dataset including 30 features. This appendix provides a detailed discussion of these experiments and their outcomes.
Real-World Datasets
To demonstrate the robustness and real-world applicability of our approach, we conducted experiments on real-world datasets: the Charlottesville City dataset, the Albemarle County dataset, Heart Disease and Breast Cancer Wisconsin.
Charlottesville City Dataset
The Charlottesville City dataset is part of the synthetic U.S. population, as described in [Chen et al., 2021] and [Barrett et al., 2009]. This dataset consists of approximately 33,000 individuals and around 5,600 activity locations visited by these individuals. The locations represent places where individuals perform various activities, providing insights into human mobility patterns and social interactions within the city.
Albemarle County Dataset
The Albemarle County dataset is another real-world dataset used in our experiments. This dataset is significantly larger than the Charlottesville City dataset, comprising about 74,000 individuals. The increased sample size allows us to evaluate the scalability and performance of our approach when applied to larger, more complex datasets.
The Albemarle County dataset contains information about individuals’ activities and the locations they visit, similar to the Charlottesville City dataset. This dataset provides a comprehensive representation of human mobility patterns and social interactions within the county.
By using these real-world datasets, we aim to validate the effectiveness and practicality of our methodology in real-life scenarios. The diverse characteristics of these datasets, such as the number of individuals and activity locations, enable us to assess the robustness and generalizability of our approach.
In our experiments, we applied our methodology to all datasets and compared the results to those obtained from the synthetic datasets. The consistency of results across these real-world datasets further reinforces the reliability and potential of our approach for real-world applications.
@@
We carefully created a synthetic dataset that mimics the properties of real-world datasets, striking a balance between realism and controlled variability. This dataset consists of 1000 uniformly distributed data points in a 2D space, with a range similar to the real datasets we analyzed.
The primary motivation behind this synthetic dataset is to provide a sandbox environment free from the unpredictable noise and anomalies of real-world data. This controlled setting is pivotal in understanding the core effects of differential privacy mechanisms, isolating them from external confounding factors. The dataset is a foundational tool in our experiments, allowing us to draw comparisons and validate our methodologies before applying them to more complex, real-world scenarios.
@@ D.3.2 Experimental Results
We observed consistent trends and patterns in all our data sets, including Charlottesville city, Albemarle county, and notably the Heart Disease dataset and the Breast Cancer dataset from the UCI Machine Learning Repository. As the value of $\epsilon$ increased to prioritize accuracy, we observed a gradual reduction in privacy protection. However, in line with our hypothesis, the impact of the epsilon budget on the explainability of our outcomes remained minimal. This consistency held true across all dimensions of the Heart Disease dataset, reinforcing the robustness of our findings across diverse data sources and attributes.
@@ D.3.3 Performance Evaluation:
For each $\epsilon$ value, we conducted 100 different runs for each dataset. The average results were consistent with our findings across all datasets. It’s essential to note that these multiple invocations were solely for performance evaluation. In real-world applications, invoking private algorithms multiple times could degrade the privacy guarantee.
@@
Despite the distinct scales between our different datasets, we observed consistent patterns in the contrastive explanations. Specifically, as illustrated in all Figures - (b), contrastive explanations remained largely unaffected by variations in the $\epsilon$ budget. This consistency further reinforces our hypothesis that the epsilon budget has a negligible influence on the explainability of our outcomes, even when applied to datasets of different scales.
@@ D.3.4 Conclusion:
The extended experiments on all our datasets further validate our approach’s efficacy. The balance between privacy and utility, the robustness of contrastive explanations, and the negligible impact of $\epsilon$ on explainability were consistent across datasets. These findings underscore the potential of our method for diverse real-world applications.